# Mildly Accurate Computationally Differentially Private Inner Product Protocols Imply Oblivious Transfer

### v1

@@ Abstract
In distributed differential privacy, multiple parties collaboratively analyze their combined data while protecting the privacy of each party’s data from the eyes of the others. Interestingly, for certain fundamental two-party functions like inner product and Hamming distance, the accuracy of distributed solutions significantly lags behind what can be achieved in the centralized model. However, under computational differential privacy, these limitations can be circumvented using oblivious transfer via secure multi-party computation. Yet, no results show that oblivious transfer is indeed necessary for accurately estimating a non-Boolean functionality. In particular, for the inner-product functionality, it was previously unknown whether oblivious transfer is necessary even for the best possible constant additive error.
@@ 1 Introduction
Much of the research in differential privacy operates within the central model. This model assumes that data owners are willing to share their data with a central server, which applies a differentially private mechanism and releases the results to the (potentially adversarial) world. However, in many scenarios, there is no such trusted server, and the data is distributed among several parties who wish to perform a joint computation while preserving each party’s data from the eyes of the others. But constructing useful multiparty DP protocols is harder than constructing them in the centralized model because now we need to develop a ${\sf DP}$ interactive protocol, which means that not only the output should be a differentially private function of the data (as in the centralized model), but also each party’s view in the interaction should be a differentially private function of the other parties’ data ([DN04, BNO08]).
@@
Perhaps the most fundamental and natural example that best illustrates the hardness of distributed ${\sf DP}$ is the inner-product functionality of two binary datasets. Namely, two parties $\mathsf{A}$ and $\mathsf{B}$ hold datasets $x\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ (respectively), and wish to estimate the inner product over the integers $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}\cdot y_{i}$ (the correlation between the datasets) using a ${\sf DP}$ protocol. We know that there exists a simple randomized response $\varepsilon$-${\sf DP}$ protocol that estimates the inner-product up to additive error $O_{\varepsilon}(\sqrt{n})$ ([McGregorMPRTV10, MPRV09]), and this is essentially the best error rate we can achieve. This was shown by [McGregorMPRTV10, HaitnerMST22] who showed that any $\varepsilon$-${\sf DP}$ protocol for the inner-product, even when allowing any $\delta\in o(1/n)$, must have an additive error of $\Omega_{\varepsilon}(\sqrt{n})$. Note that this is in complete contrast to the centralized model, where the simple $\varepsilon$-${\sf DP}$ Laplace mechanism $\mathsf{M}(x,y)=\langle x,y\rangle+{\rm Lap}(2/\varepsilon)$ has only additive error $O_{\varepsilon}(1)$.
@@
To overcome such accuracy gaps between the models, [BNO08, MPRV09] considered protocols that only guarantee a computational analog of differential privacy (denote by ${\sf CDP}$), which is a relaxation of Section 1 where Equation 1 only holds for computationally efficient (ppt) algorithms $\mathsf{D}$, and not necessarily for any algorithm $\mathsf{D}$. This relaxation is very powerful. For example, if we assume the existence of the cryptographic primitive oblivious transfer (in short, ${\sf OT}$, [Rabin81]), then in any distributed setting (even for more than two parties), we can achieve the accuracy of a centralize ${\sf DP}$ mechanism by emulating it using secure multiparty computation ([BNO08, DKMMN06]). In particular, using ${\sf OT}$, two parties can securely emulate the (centralized) Laplace mechanism to achieve an $\varepsilon$-${\sf CDP}$ protocol with additive error $O_{\varepsilon}(1)$. This leads to the following fundamental question:
@@
[HaitnerMST22] only made partial progress on this matter, showing that any ${\sf CDP}$ protocol with non-trivial accuracy guarantee of $o_{\varepsilon}(\sqrt{n})$ must use a cryptographic primitive that is at least as strong as key agreement (in short, ${\sf KA}$). But although ${\sf KA}$ and ${\sf OT}$ are both public-key cryptography assumptions, it is known that ${\sf KA}$ is strictly weaker than ${\sf OT}$ (at least under black-box reductions [GertnerKMRV00]). In fact, even for ${\sf CDP}$ protocols that estimate the inner-product with very high accuracy of $O_{\varepsilon}(1)$, it was previously unknown whether ${\sf OT}$ is indeed necessary.
@@ 1.1 Our Result
In this work, we prove the first equivalence between ${\sf OT}$ and ${\sf CDP}$ two-party protocols that estimate the inner-product with small additive error, and in particular, additive error up to $O_{\varepsilon}(n^{1/6})$.
@@
An $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ two-party protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $n^{1/6}/(c\cdot e^{c\varepsilon})$ with probability $0.999$ (for some universal constant $c>0$), can be used to construct an ${\sf OT}$ protocol. [^1] We remark that Theorem 1.1 is proven via a uniform and black-box reduction. Thus, although the security of ${\sf CDP}$ and ${\sf OT}$ is commonly defined with respect to non-uniform adversaries, Theorem 1.1 also implies a reduction from ${\sf OT}$ that is secure against uniform adversaries to such mildly accurate inner-product protocols that are ${\sf CDP}$ with respect to uniform distinguishers.
In particular, Theorem 1.1 implies that without assuming the existence of ${\sf OT}$, it is impossible to construct a ${\sf CDP}$ protocol for the inner-product with a constant, or even sub-polynomial, additive error.
As a direct corollary of Theorem 1.1, we obtain that a protocol with a small polynomial additive error is equivalent to a protocol with a constant error. That is because using such a mildly accurate protocol, we can construct by Theorem 1.1 an oblivious transfer, and then use it to perform a secure two-party computation of the Laplace mechanism.
@@
There exists a universal constant $c>0$ such that the following holds. Assume that an $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ two-party protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $n^{1/6}/(c\cdot e^{c\varepsilon})$ with probability $0.999$ exists. Then there exists an $\mathopen{}\mathclose{{}\left(\varepsilon,\operatorname{neg}(n)}\right)$-${\sf CDP}$ protocol that estimates the inner product up to an additive error $20/\varepsilon$ with probability $0.999$.
We remark that Theorem 1.1 is the first equivalence result between ${\sf OT}$ to any ${\sf CDP}$ protocol that estimates a natural, non-Boolean functionality. Prior works only provide such equivalences for ${\sf CDP}$ protocols that estimate the (Boolean) XOR functionality ([GKMPS16, HMSS19]), which is less interesting in the context of ${\sf DP}$, as even in the centralized model, any ${\sf DP}$ algorithm for estimating XOR must incur an error close to one-half. In contrast, the inner product has a significantly larger gap between the accuracy achievable in two-party ${\sf DP}$ and in ${\sf CDP}$.
A proof overview of Theorem 1.1 is given in Section 2.
@@ 1.2 Open Questions
In this paper, we show that a $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $O_{\varepsilon}(n^{1/6})$ is equivalent to oblivious transfer (${\sf OT}$). For the inner product functionality, it still left open to close the gap up to any non-trivial accuracy of $O_{\varepsilon}(\sqrt{n})$. [HaitnerMST22] closed this gap w.r.t. a weaker notion of ${\sf CDP}$ against external observer [^2] A protocol is ${\sf CDP}$ against external observer, if the transcript of the execution is a ${\sf CDP}$ function of the input datasets, but not necessarily the views of the parties as in the standard notion. by showing that key agreement (${\sf KA}$) is necessary and sufficient for any non-trivial accuracy guarantee, which in particular implies that ${\sf KA}$ is also necessary for the standard notion of ${\sf CDP}$. But since ${\sf KA}$ is strictly weaker than ${\sf OT}$ (under block-box reductions), we still do not know what is the right answer for additive errors in the range between $n^{1/6}$ and $\sqrt{n}$.
@@ 1.3 Paper Organization
[…]
@@ 2 Proof Overview
    Next, $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=0}\right\}$.
@@
The first observation made by [HaitnerMST22] is that, by linearity of the inner product, $\langle x_{-r},y_{-r}\rangle+\langle x_{r},y_{r}\rangle=\langle x,y\rangle$, and thus $\mathopen{}\mathclose{{}\left|{\rm out}_{A}-{\rm out}_{B}}\right|=\mathopen{}\mathclose{{}\left|z-\langle x,y\rangle}\right|$, which is small assuming $\Pi$ is accurate. Moreover, it was shown that this protocol can be amplified into a full-fledged ${\sf KA}$ protocol if any efficient adversary $\mathsf{E}$ cannot approximate ${\rm out}_{A}$ within small additive distance, given the transcript of the weak ${\sf KA}$ protocol. The main technical part in [HaitnerMST22] is to show that this is indeed the case in the above protocol.
In this work, we are interested in constructing a stronger primitive — that is, oblivious transfer (${\sf OT}$). In contrast to a ${\sf KA}$ protocol in which security must hold against external observers that only see the transcript, in ${\sf OT}$, the security needs to hold against an adversary that has access to the full view of one of the parties. Our first observation is that the original ${\sf KA}$ protocol of [HaitnerMST22], as is, cannot be used to construct any form of even “weak” oblivious transfer. In more details, it is possible to show that for some carefully chosen ${\sf CDP}$ and accurate protocol $\Pi$, the joint view of the parties in the above protocol can be simulated using a trivial protocol, without using $\Pi$ at all. [^3] Here a protocol $\Pi^{\prime}$ simulates $\Pi$ if $\Pi^{\prime}$ generates the joint view of the parties $\mathsf{A}$, $\mathsf{B}$ in $\Pi$, without revealing any other information to the parties. We say that the simulation $\Pi^{\prime}$ is trivial if it does not rely on any nontrivial protocol or cryptographic assumptions. Importantly, the simulation may reveal new information (that is already known for the internal parties $\mathsf{A}$ and $\mathsf{B}$) to an external observer, and thus it doesn’t imply that the protocol is not a ${\sf KA}$ protocol. See Appendix B for a concrete example. Since there is no ${\sf OT}$ without computational assumptions, this leads to the conclusion that the ${\sf KA}$ protocol of [HaitnerMST22] could not be used as a subroutine (in a black-box manner) to construct ${\sf OT}$. Looking ahead, our solution is to inject carefully placed noise to the interaction, in a manner that cannot be simulated by the parties without revealing private information. As we will see shortly, this enables us to bypass the limitations of the original construction of [HaitnerMST22] and to achieve ${\sf OT}$.
@@
Before describing our construction, let us explain the properties which we need the construction to fulfill. Similarly to [HaitnerMST22], we will start with constructing a weak ${\sf OT}$ protocol, and then we will show how to amplify it into a full-fledged ${\sf OT}$ protocol. By the work of [Wullschleger09], to achieve OT, it is sufficient to construct a weak erasure channel (WEC). [^4] A weak erasure channel is a weak version of Rabin’s ${\sf OT}$ [Rabin81], which is known to be equivalent to (${2\choose 1}$-${\sf OT}$. In this work, we construct a protocol with a weaker security guarantee, which we call approximate weak erasure channel (AWEC), and show that such a protocol also suffices for constructing ${\sf OT}$. Informally, an $(\ell,\alpha,p,q)$-AWEC is a protocol between two parties $\mathsf{A}$ and $\mathsf{B}$ that have no input, in which $\mathsf{A}$ outputs a number $O_{\mathsf{A}}\in[-n,n]$, and $\mathsf{B}$ either outputs a number $O_{\mathsf{B}}\in[-n,n]$, or an erasure symbol $O_{\mathsf{B}}=\bot$. We additionally require that:
@@
  • When an erasure occurs, $\mathsf{B}$ cannot predict the value of $O_{A}$. That is, let $V_{B}$ be the view of $\mathsf{B}$ in the protocol. Then we require that for any efficient algorithm $\widehat{\mathsf{B}}$ it holds that
@@
  • Lastly, we require that $\mathsf{A}$ cannot say when an erasure occurs. That is, let $V_{A}$ be the view of $\mathsf{A}$ in the protocol. Then we require that for any efficient algorithm $\widehat{\mathsf{A}}$ it holds that
    $\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A})=1\mid O_{B}\neq\perp}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A})=1\mid O_{B}=\perp}\right]}\right|\leq p$
@@
Operation: 1. $\mathsf{A}$ samples $x\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$, and $\mathsf{B}$ samples $y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$. 2. $\mathsf{A}$ and $\mathsf{B}$ interact according to $\Pi$ using inputs $x$ and $y$ resp., to get output $z\approx\langle x,y\rangle\pm\ell$. 3. $\mathsf{A}$ samples $r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and sends $(r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\})$ to $\mathsf{B}$. 4. $\mathsf{B}$ samples a random bit $b\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and acts as follows: (a) If $b=0$, it sends $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$ to $\mathsf{A}$ and outputs $o_{B}=z-\langle x_{r},y_{r}\rangle$. (b) Otherwise ($b=1$), it performs the following steps: Samples $k$ uniformly random indices $i_{1},\ldots,i_{k}\leftarrow[n]$, where $n^{1/3}\gg k\gg e^{\varepsilon}\cdot\ell^{2}$. Compute $\widetilde{y}=(\widetilde{y}_{1},\ldots,\widetilde{y}_{n})$ where $\widetilde{y}_{i}\leftarrow\begin{cases}{\cal U}_{\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}}&i\in\mathopen{}\mathclose{{}\left\{i_{1},\ldots,i_{k}}\right\}\\ y_{i}&\text{otherwise}\end{cases}$, Send $\widetilde{y}_{-r}=\mathopen{}\mathclose{{}\left\{\widetilde{y}_{i}\colon r_{i}=0}\right\}$ to $\mathsf{A}$, and Output $o_{B}=\perp$. 5. Denote by $\widehat{y}_{-r}$ the message $\mathsf{A}$ received from $\mathsf{B}$. Then $\mathsf{A}$ outputs $o_{A}=\langle x_{-r},\widehat{y}_{-r}\rangle$.

### v2

# Computationally Differentially Private Inner-Product Protocols Imply Oblivious Transfer

@@ Abstract
In distributed differential privacy, multiple parties collaborate to analyze their combined data while each party protects the confidentiality of its data from the others. Interestingly, for certain fundamental two-party functions, such as the inner product and Hamming distance, the accuracy of distributed solutions significantly lags behind what can be achieved in the centralized model. For computational differential privacy, however, these limitations can be circumvented using oblivious transfer (used to implement secure multi-party computation). Yet, no results show that oblivious transfer is indeed necessary for accurately estimating a non-Boolean functionality. In particular, for the inner-product functionality, it was previously unknown whether oblivious transfer is necessary even for the best possible constant additive error.
@@ 1 Introduction
Most research in differential privacy regards the centralized model: users share their data with a central server that applies a differentially private mechanism and releases the results to the (potentially adversarial) world. In many scenarios, however, there is no such trusted server, and the data is distributed among several parties who wish to perform a joint computation while preserving the privacy of each party’s data from the eyes of the others. Constructing multiparty protocols that realize a functionality while preserving ${\sf DP}$ is typically more challenging than creating them in the centralized model; a party’s output should not only be a differentially private given the other parties’ inputs and outputs (as in the centralized model), but also given the other parties’ view in the protocol ([DN04, BNO08]).
@@
A fundamental example that illustrates the hardness of distributed ${\sf DP}$ is the inner-product functionality of two binary datasets: parties $\mathsf{A}$ and $\mathsf{B}$ hold datasets $x\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ (respectively), and wish to estimate the inner product over the integers $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}\cdot y_{i}$ (the correlation between the datasets) using a ${\sf DP}$ protocol. There exists a protocol (specifically, the so-called randomized response) that is $\varepsilon$-${\sf DP}$ and estimates the inner-product up to additive error $O_{\varepsilon}(\sqrt{n})$ ([McGregorMPRTV10, MPRV09]), and this essentially the best error rate we can achieve while maintaining $\varepsilon$-${\sf DP}$: [McGregorMPRTV10, HaitnerMST22] showed that any $\varepsilon$-${\sf DP}$ protocol for the inner-product, even when allowing any $\delta\in o(1/n)$, must have an additive error of $\Omega_{\varepsilon}(\sqrt{n})$. This is in contrast to the centralized model, where the simple $\varepsilon$-${\sf DP}$ Laplace mechanism $\mathsf{M}(x,y)=\langle x,y\rangle+{\rm Lap}(2/\varepsilon)$ has only additive error $O_{\varepsilon}(1)$.
@@
To achieve better accuracy, [BNO08, MPRV09] considered protocols that only guarantee computational differential privacy (${\sf CDP}$): a computational analog of ${\sf DP}$ where Section 1 is relaxed so that Equation 1 only holds for computationally efficient (ppt) algorithms $\mathsf{D}$, and not necessarily for any algorithm $\mathsf{D}$. This relaxation is very powerful. For example, if we assume the existence of the cryptographic primitive oblivious transfer (in short, ${\sf OT}$, [Rabin81]), then in any distributed setting (even for more than two parties), we can achieve the accuracy of a centralize ${\sf DP}$ mechanism by emulating it using secure multiparty computation ([BNO08, DKMMN06]). In particular, using ${\sf OT}$, two parties can securely emulate the (centralized) Laplace mechanism to achieve an $\varepsilon$-${\sf CDP}$ protocol with additive error $O_{\varepsilon}(1)$. This leads to the following fundamental question:
@@
[…]
*HaitnerMST22 made partial progress on this matter, showing that any $(\varepsilon,\frac{1}{n^{2}})$-${\sf CDP}$ protocol with non-trivial accuracy guarantee of $o_{\varepsilon}(\sqrt{n})$ must use a cryptographic primitive that is at least as strong as key agreement (in short, ${\sf KA}$). However, although ${\sf KA}$ and ${\sf OT}$ are both public-key cryptography primitives, ${\sf KA}$ is believed to be a weaker primitive ([GertnerKMRV00]) and does not suffice for secure MPC. Even for ${\sf CDP}$ protocols that estimate the inner-product with very high accuracy of $O_{\varepsilon}(1)$, it was previously unknown whether ${\sf OT}$ is indeed necessary.
@@ 1.1 Our Result
We prove the first equivalence between ${\sf OT}$ and ${\sf CDP}$ two-party protocols that estimate the inner-product with small additive error, and in particular, additive error up to $O_{\varepsilon}(n^{1/6})$.
@@
An $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ two-party protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $n^{1/6}/(c\cdot e^{c\varepsilon})$ with probability $0.999$ (for some universal constant $c>0$), can be used to construct an ${\sf OT}$ protocol. [^1] Theorem 1.1 is proven via a uniform and black-box reduction. Thus, although the security of ${\sf CDP}$ and ${\sf OT}$ is commonly defined with respect to non-uniform adversaries, Theorem 1.1 also implies a reduction from ${\sf OT}$ that is secure against uniform adversaries to such mildly accurate inner-product protocols that are ${\sf CDP}$ with respect to uniform distinguishers.
In particular, Theorem 1.1 implies that without assuming the existence of ${\sf OT}$, it is impossible to construct a ${\sf CDP}$ protocol for the inner-product with a constant, or even sub-polynomial, additive error. It follows that a protocol with a small polynomial additive error is equivalent to a protocol with a constant error. That is because using such a mildly accurate protocol, we can construct by Theorem 1.1 an oblivious transfer, and then use it to perform a secure two-party computation of the Laplace mechanism.
@@
There exists a universal constant $c>0$ such that the following holds: assume that an $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ two-party protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $n^{1/6}/(c\cdot e^{c\varepsilon})$ with probability $0.999$ exists, then there exists an $\mathopen{}\mathclose{{}\left(\varepsilon,\operatorname{neg}(n)}\right)$-${\sf CDP}$ protocol that estimates the inner product up to an additive error $20/\varepsilon$ with probability $0.999$.
Theorem 1.1 is the first equivalence result between ${\sf OT}$ to any ${\sf CDP}$ protocol that estimates a natural, non-Boolean functionality. Prior works only provide such equivalences for ${\sf CDP}$ protocols that estimate the (Boolean) XOR functionality ([GKMPS16, HMSS19]), a less interesting functionality in the context of ${\sf DP}$, as even in the centralized model, any ${\sf DP}$ algorithm for estimating XOR must incur an error close to one-half. In contrast, the inner product has a significantly larger gap between the accuracy achievable in two-party ${\sf DP}$ and in ${\sf CDP}$. A proof overview of Theorem 1.1 is given in Section 2.
@@ 1.2 Open Questions
In this paper, we show that a $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $O_{\varepsilon}(n^{1/6})$ is equivalent to oblivious transfer (${\sf OT}$). For the inner-product functionality, it is still left open to close the gap up to any non-trivial accuracy of $O_{\varepsilon}(\sqrt{n})$. [HaitnerMST22] closed this gap w.r.t. a weaker notion of ${\sf CDP}$ against external observer [^2] A protocol is ${\sf CDP}$ against external observer, if the transcript of the execution is a ${\sf CDP}$ function of the input datasets, but not necessarily the views of the parties as in the standard notion. by showing that key agreement (${\sf KA}$) is necessary and sufficient for any non-trivial accuracy guarantee, which in particular implies that ${\sf KA}$ is also necessary for the standard notion of ${\sf CDP}$. But since ${\sf KA}$ is strictly weaker than ${\sf OT}$ (under block-box reductions), we still do not know what is the right answer for additive errors in the range between $n^{1/6}$ and $\sqrt{n}$.
@@ 1.3 Paper Organization
In Section 2, we give a high-level proof of Theorem 1.1, and in Section 3, we discuss additional related works. Notations, definitions, and general statements used throughout the paper are given in Section 4. Our main formal theorems are stated in Section 5, and the oblivious transfer protocol together with its security proof are given in Section 6. In Section 3 we discuss additional related works. The proof of our main result relies on a technical tool that is proven in Appendix B. Other missing proofs appear in Appendix C.
@@ 2 Proof Overview
    Next, $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$.
@@
The first observation made by [HaitnerMST22] is that, by linearity of the inner product, $\langle x_{-r},y_{-r}\rangle+\langle x_{r},y_{r}\rangle=\langle x,y\rangle$, and thus $\mathopen{}\mathclose{{}\left|{\rm out}_{A}-{\rm out}_{B}}\right|=\mathopen{}\mathclose{{}\left|z-\langle x,y\rangle}\right|$, which is small assuming $\Pi$ is accurate. Moreover, it was shown that this protocol can be amplified into a full-fledged ${\sf KA}$ protocol if any efficient adversary $\mathsf{E}$ cannot approximate ${\rm out}_{A}$ within a small additive distance, given the transcript of the weak ${\sf KA}$ protocol. The main technical part in [HaitnerMST22] is to show that this is indeed the case in the above protocol.
In this work, we are interested in constructing a stronger primitive, namely oblivious transfer (${\sf OT}$). In contrast to a ${\sf KA}$ protocol in which security must hold against external observers that only see the transcript, in ${\sf OT}$, the security needs to hold against an adversary that has access to the full view of one of the parties. Our first observation is that the original ${\sf KA}$ protocol of [HaitnerMST22], as is, cannot be used to construct any form of even “weak” oblivious transfer. In more details, it is possible to show that for some carefully chosen ${\sf CDP}$ and accurate protocol $\Pi$, the joint view of the parties in the above protocol can be simulated using a trivial protocol, without using $\Pi$ at all. [^3] Here a protocol $\Pi^{\prime}$ simulates $\Pi$ if $\Pi^{\prime}$ generates the joint view of the parties $\mathsf{A}$, $\mathsf{B}$ in $\Pi$, without revealing any other information to the parties. We say that the simulation $\Pi^{\prime}$ is trivial if it does not rely on any nontrivial protocol or cryptographic assumptions. Importantly, the simulation may reveal new information (that is already known for the internal parties $\mathsf{A}$ and $\mathsf{B}$) to an external observer, and thus it doesn’t imply that the protocol is not a ${\sf KA}$ protocol. See Appendix A for a concrete example. Since there is no ${\sf OT}$ without computational assumptions, this leads to the conclusion that the ${\sf KA}$ protocol of [HaitnerMST22] could not be used as a subroutine (in a black-box manner) to construct ${\sf OT}$. Looking ahead, our solution is to inject carefully placed noise to the interaction, in a manner that the parties cannot simulate without revealing private information. As we will see shortly, this enables us to bypass the limitations of the original construction of [HaitnerMST22] and to achieve ${\sf OT}$.
@@
Before describing our construction, let us explain the properties that we need the construction to fulfill. Similarly to [HaitnerMST22], we will start with constructing a weak ${\sf OT}$ protocol, and then we will show how to amplify it into a full-fledged ${\sf OT}$ protocol. By the work of [Wullschleger09], to achieve OT, it is sufficient to construct a weak erasure channel (WEC). [^4] A weak erasure channel is a weak version of Rabin’s ${\sf OT}$ [Rabin81], which is known to be equivalent to (${2\choose 1}$-${\sf OT}$. In this work, we construct a protocol with a weaker security guarantee, which we call approximate weak erasure channel (AWEC), and show that such a protocol also suffices for constructing ${\sf OT}$. Informally, an $(\ell,\alpha,p,q)$-AWEC is a no-input protocol between $\mathsf{A}$ and $\mathsf{B}$, party $\mathsf{A}$ outputs a number $O_{\mathsf{A}}\in[-n,n]$, and party $\mathsf{B}$ either outputs a number $O_{\mathsf{B}}\in[-n,n]$, or an erasure symbol $O_{\mathsf{B}}=\bot$. We additionally require that:
@@
  • When an erasure occurs, $\mathsf{B}$ cannot predict the value of $O_{A}$: let $V_{B}$ be the view of $\mathsf{B}$ in the protocol, then for any efficient $\widehat{\mathsf{B}}$
@@
  • Lastly, we require that $\mathsf{A}$ cannot tell when an erasure occurs: let $V_{A}$ be the view of $\mathsf{A}$ in the protocol, then for any efficient $\widehat{\mathsf{A}}$
    $\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A})=1\mid O_{B}\neq\perp}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A})=1\mid O_{B}=\perp}\right]}\right|\leq p.$
@@
Operation: 1. $\mathsf{A}$ samples $x\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$, and $\mathsf{B}$ samples $y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$. 2. $\mathsf{A}$ and $\mathsf{B}$ interact according to $\Pi$ using inputs $x$ and $y$ resp., to get output $z\approx\langle x,y\rangle\pm\ell$. 3. $\mathsf{A}$ samples $r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and sends $(r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\})$ to $\mathsf{B}$. 4. $\mathsf{B}$ samples a random bit $b\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and acts as follows: (a) If $b=0$ (“non-erasure”), it sends $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$ to $\mathsf{A}$ and outputs $o_{B}=z-\langle x_{r},y_{r}\rangle$. (b) Otherwise ($b=1$, “erasure”), it performs the following steps: Samples $k$ uniformly random indices $i_{1},\ldots,i_{k}\leftarrow[n]$, where $n^{1/3}\gg k\gg e^{\varepsilon}\cdot\ell^{2}$. Compute $\widetilde{y}=(\widetilde{y}_{1},\ldots,\widetilde{y}_{n})$ where $\widetilde{y}_{i}\leftarrow\begin{cases}{\cal U}_{\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}}&i\in\mathopen{}\mathclose{{}\left\{i_{1},\ldots,i_{k}}\right\}\\ y_{i}&\text{otherwise}\end{cases}$, Send $\widetilde{y}_{-r}=\mathopen{}\mathclose{{}\left\{\widetilde{y}_{i}\colon r_{i}=0}\right\}$ to $\mathsf{A}$, and Output $o_{B}=\perp$. 5. Denote by $\widehat{y}_{-r}$ the message $\mathsf{A}$ received from $\mathsf{B}$. Then $\mathsf{A}$ outputs $o_{A}=\langle x_{-r},\widehat{y}_{-r}\rangle$.